OpenStudy (anonymous):

Write an equation of a line with the given slope and y-intercept. m = 1/4 b=-3/4

OpenStudy (zale101):

y=mx+b is the slope intercept form now replace the numbers 1/4 for m and -3/4 for b

OpenStudy (anonymous):

y=1/4x-3/4 y=3/4x+1/4 y=1/4x+3/4 y=4x-3/4

OpenStudy (anonymous):

my choices^

OpenStudy (zale101):

i told you the answer, so what is your choice?

OpenStudy (anonymous):

y=1/4+-3/4

OpenStudy (zale101):

yes what happens when negative and postive numbers come together?

OpenStudy (zale101):

when positive and negative signs come together, it forms a negative sign :)

OpenStudy (zale101):

so this y=1/4+-3/4 will become like this y=1/4-3/4

OpenStudy (zale101):

A is the answer

OpenStudy (anonymous):

ok :) thank you